OpenStudy (anonymous):

Help complete the square f(x)=-x^2-4x-8

OpenStudy (anonymous):

again start with \[f(x)=-(x+4)-8\] then \[f(x)=-(x+2)^2+k\] and find \(k\) via \(k=f(-2)\)

OpenStudy (anonymous):

sorry i meant start with \[f(x)=-(x^2+4x)-8\]

OpenStudy (anonymous):

I believe you take \[-x^2-4x-8\] and set it equal to zero. You then proceed to isolate the -x^2 - 4x. So then you have \[-x^2-4x=8\] The next thing I did was to multiply through by -1 on both sides. \[x^2+4x=-8\] Now to complete the square you have to get the x^2+4x into the form (x+c)(x+c)=(x+c)^2 To do this you divide the coefficient of the 4x by 2 and then square it. So you should get positive 4. The next step is to add this 4 to both sides of \[x^2+4x=-8\] So you should get \[x^2+4x+4=-4\] Simplifying the equation gives \[(x+2)^2=-4 \] or \[(x+2)^2+4=0\]